what would the amount of a check issued for $11283.45 paid on 7/24 with the following terms 3/10, 2/15, N 30- 5% APR
Answer #1

According the above question the discount terms is 3/10,2/5,N 30- 5%APR

If the invoice is paid with in between 3 t0 10 days from invoice issued then discount =3%of invoice.cheque amount=$10,944.9

if the invoice is paid with in between 10 to 15 days from invoice issued then discount=2%of invoice.cheque amount=$11,057.781

the invoice is paid with in between 15 to 30 days from invoice issued then discount=no discount.cheque amount=$11283.45

the invoice is paid after 30 days of invoice is issued then 5% of APR IS ADDED TO BILL

HERE CHEQUE PAID 7/24

But the when the invoice is issued is not given in the problem.Assuming the same month wnen the cheque paid.

then the cheqe amount is $11,283.45

assuming the invoice is issued begginin of the financial year i.e 4\1

then thegap between 4/1 to 7/24 is=115 days then interest =11,283.45*5%*115/365=177.75

the cheque amount is 11,461.2